I am stuck on this!

Point charge q1 = −5.00×10−6 C is on the x-axis at x = -0.400 m . Point charge q2 is on the x-axis at x = +0.6 m . Point charge q3 = +3.00×10^−6 C is at the origin.

A) What is q2 (magnitude and sign) if the net force on q3 is 6.00 N in the +x-direction?

B) What is q2 (magnitude and sign) if the net force on q3 is 6.00 N in the −x-direction?

Please help! Thanks!
Answer #1

9 = 0.4m I h = 0.6m te = kazan (-5) f23 = KA2 (i) : : t23 + tis= 6 kąaz - + kq, q 32 926 (92109) (4) (3x154 (0.432 2100) (-5*5 x 10-6 (920) (**) home - Speed ) = 6 La ville ---- 0.62 - 5 x 106 0.16 0.36 - 13.8 x 10-6 b = 2-2 X104 2-7 8104 0.16 = 2-2